OpenStudy (anonymous):

[(radx+4)+(radx-1)]^2

OpenStudy (anonymous):

\[(\sqrt{x+4} - \sqrt{x-4})^{2}\]

OpenStudy (anonymous):

\[(a-b)^{2} = a^{2} -2ab + b^{2}\] a = \[\sqrt{x + 4}\] b = \[\sqrt{x-1}\]

OpenStudy (anonymous):

well at the top it says (radx-1) and below it says (radx-4) which one is it really?

OpenStudy (anonymous):

top

OpenStudy (anonymous):

than just do what I said instead b = \[\sqrt{x-4}\]

OpenStudy (anonymous):

okay, thanks!